I am not sure how to work this sample question out. I don't know where to start
Suppose you want to cover the backyard with decorative rock and plant some trees as the first phase of the project. You need 30 tons of rock to cover the area. If each ton cost $60 and each tree is $84, what is the maximum number of trees you can buy with a budget for rock and trees of $2,500? Write an inequality that illustrates the problem and solve.
Would 5 trees be a solution to the inequality in part b?

Answer by Earlsdon(6287)

Let T = the number of trees.
The cost of the rock will be:
(30 tons)*($60 per ton) = $1800
We can write the inequality:
$1800+($84)*T <= $2500 Subtract $1800 from both sides.
($84)*T <= $700 Divide both sides by $84.
T <= 8.3333 but you can't buy a fraction of a tree, so...
T <= 8
You could buy a maximum of 8 trees.
Yes, 5 trees would be a solution because 5 <= 8